K is a point located between J and L. If Jk=2x+5 and KL=5x+3 and JL=71,find the measure of JK and KL
Answer by cleomenius(959)

The line is 71.
The line segments will total 71.
2x + 5 + 5x + 3 = 71
7x + 8 = 71
7x = 63
x = 9
2(9) + 5 + 5(9) + 3 = 71
18 + 5 + 45 + 3 = 71
Cleomenius.